40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 293 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  2798   Tue Apr 13 12:49:35 2010 josephbUpdateComputersY end simulated plant progress

Quote:

Currently, the y end plant is yep.mdl.  In order to compile it properly (for the moment at least) requires running the normal makefile, then commenting out the line in the makefile which does the parsing of the mdl, and rerunning after modifying the /cds/advLigo/src/fe/yep/yep.c file.

The modifications to the yep.c file are to change the six lines that look like:

"plant_mux[0] = plant_gndx"  into lines that look like "plant_mux[0] = plant_delayx".  You also have to add initialization of the plant_delayx type variables to zero in the if(feInt) section, near where plant_gndx is set to zero.

This is necessary to get the position feedback within the plant model to work properly.

 

#NOTE by Koji

CAUTION:
This entry means that Makefile was modified not to parse the mdl file.
This affects making any of the models on megatron.

 To prevent this confusion in the future, at Koji's suggestion I've created a Makefile.no_parse_mdl in /home/controls/cds/advLIGO on megatron.  The normal makefile is the original one (with correct parsing now).  So the correct procedure is:

1) "make yep"

2) Modify yep.c code

3) "make -f Makefile.no_parse_mdl yep"

  2797   Tue Apr 13 12:39:51 2010 Aidan, MottSummaryGreen LockingTemperature sweep of the Innolight: df/dT ~ 3.3GHz/K

Please put those numbers onto wiki somewhere at the green page or laser characterization page.

Quote:

Quote:

The beams from the Innolight and Lightwave NPROs were both incident on a 1GHZ New Focus PD. Mott and I swept the temperature of the Lightwave and tracked the change in frequency of the beatnote between the two. The Innolight temperature was set to 39.61C although the actual temperature was reported to be 39.62C.

Freq. vs temperature is plotted below in the attached PDF. The slope is 2.8GHz/K.

The data is in the attached MATLAB file.

 Same thing for the Innolight Mephisto.

Not unexpected values with dn/dT around 11E-6 K^-1 and coefficient of thermal expansion = 8E-6 K^-1 and a laser resonator length of order 10cm.

 

  2796   Mon Apr 12 22:51:31 2010 KojiUpdateSUSITMX installed and aligned

Koji

ITMX was aligned with regard to the 40m green oplev.
Now both cavities are aligned.

Next thing we are going to do is to remove PRM and SRM towers.

As well as the oplev construction for ITMs.

We anticipate the drift of the stack. So we need to revisit the alignment again.

Some tools and the level gauge were removed from the table.

Picture of the ETMX - reflection from the ITMX is hitting the mirror and Jamie's windmill.

 IMG_2381.jpg

 


0. The suspension tower had been placed on the table close to the door.

1. Brought the OSEMs from the clean room. Connected the satellite box to the ITMX suspension.

2. Went into the chamber. Leveled the table.

3. Released the mirror from the clamp. Put and adjust the OSEMs.

- Note that the side OSEM is located to the south side of the tower
  so that we can still touch it after the placement of the TT suspension at the north side of the SOS tower.

4. Clamped the mirror. Moved the SOS tower according to the CAD layout.

5. Leveled the table again.

6. Released the mirror again and adjusted the OSEMs.

7. Turned on the end green laser pointer.

- The spot was slightly upside and left of the mirror. Adjusted it so that the spot is at the center.

8. Align ITMX in Pitch

- The spot was hitting the tube. Moved the pitch bias such that the beam get horizontal.

9. Align ITMX in Yaw

- Moved the SOS tower such that the approximate spot is on the ETMX. If I hit the right spot I could see the tube get grown green because of the huge scatter.

10. Adjusted the OSEMs again and check the alignment again. Repeated this process 2~3 times.

- Bias values at  the end of the work: Pitch 0.7800 / Yaw 0.270

11. Close up the chamber

- Remove the level gauge. Some of the screws are still in the Al ship in the chamber.

- Close the light door.

  2795   Mon Apr 12 22:44:30 2010 KojiUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Data looks perfect ... but the fitting was wrong.

Vn = Vdn + Z * sqrt( 2 e Idc ) ==> WRONG!!!

Dark noise and shot noise are not correlated. You need to take a quadratic sum!!!

Vn^2 = Vdn^2 + Z^2 *(2 e Idc)

And I was confused whether you need 2 in the sqrt, or not. Can you explain it?
Note that you are looking at the raw RF output of the PD and not using the demodulated output... 

Also you should be able to fit Vdn. You should put your dark noise measurement at 10nA or 100nA and then make the fitting.

Quote:

 Here's another measurement of the noise of the REFL11 PD.

This time I made the fit constraining the Dark Noise. I realized that it didn't make much sense leaving it as a free coefficient: the dark noise is what it is.

2010-04-09_REFL11NoiseMeasurements.png

Result: the transimpedance of REFL11at 11 MHz is about 4000 Ohm.

Note:
This time, more properly, I refer to the transimpedance as the ratio between Vout @11Mhz / Photocurrent. In past entries I improperly called transimpedance the impedance of the circuit which resonates with the photodiode.

 

  2794   Mon Apr 12 20:48:51 2010 Aidan, MottSummaryGreen LockingTemperature sweep of the Innolight: df/dT ~ 3.3GHz/K

Quote:

The beams from the Innolight and Lightwave NPROs were both incident on a 1GHZ New Focus PD. Mott and I swept the temperature of the Lightwave and tracked the change in frequency of the beatnote between the two. The Innolight temperature was set to 39.61C although the actual temperature was reported to be 39.62C.

Freq. vs temperature is plotted below in the attached PDF. The slope is 2.8GHz/K.

The data is in the attached MATLAB file.

 Same thing for the Innolight Mephisto.

Not unexpected values with dn/dT around 11E-6 K^-1 and coefficient of thermal expansion = 8E-6 K^-1 and a laser resonator length of order 10cm.

Attachment 1: Innolight_temp_sweep.pdf
Innolight_temp_sweep.pdf
Attachment 2: Innolight_Temp.m
% plot the data from the Innolight Temperature sweep

% Innolight temperature

InnTemp = [0.60
    .59
    .56
    .52
    .65] + 39;

... 25 more lines ...
  2793   Mon Apr 12 19:50:30 2010 AidanSummaryGreen LockingTemperature sweep of the Lightwave: df/dT = 2.8GHz/K

The beams from the Innolight and Lightwave NPROs were both incident on a 1GHZ New Focus PD. Mott and I swept the temperature of the Lightwave and tracked the change in frequency of the beatnote between the two. The Innolight temperature was set to 39.61C although the actual temperature was reported to be 39.62C.

Freq. vs temperature is plotted below in the attached PDF. The slope is 2.8GHz/K.

The data is in the attached MATLAB file.

Attachment 1: LightWave_temp_sweep.pdf
LightWave_temp_sweep.pdf
Attachment 2: LightWave_Temp.m
% plot the data from the Lightwave Temperature sweep

% Lightwave temperature

LWTemp = [0.2744
    0.2753
    .2767
    .2780
    .2794
    .2808
... 67 more lines ...
  2792   Mon Apr 12 17:48:32 2010 AidanUpdateComputer Scripts / Programselog restarted

 The elog crashed when I was uploading a photo just now. I logged into nodus and restarted it.

  2791   Mon Apr 12 17:37:52 2010 josephbUpdateComputersY end simulated plant progress

Currently, the y end plant is yep.mdl.  In order to compile it properly (for the moment at least) requires running the normal makefile, then commenting out the line in the makefile which does the parsing of the mdl, and rerunning after modifying the /cds/advLigo/src/fe/yep/yep.c file.

The modifications to the yep.c file are to change the six lines that look like:

"plant_mux[0] = plant_gndx"  into lines that look like "plant_mux[0] = plant_delayx".  You also have to add initialization of the plant_delayx type variables to zero in the if(feInt) section, near where plant_gndx is set to zero.

This is necessary to get the position feedback within the plant model to work properly.

 

#NOTE by Koji

CAUTION:
This entry means that Makefile was modified not to parse the mdl file.
This affects making any of the models on megatron.

Attachment 1: YEP.png
YEP.png
Attachment 2: YEP_PLANT.png
YEP_PLANT.png
  2790   Mon Apr 12 17:09:30 2010 AlbertoUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Quote:

Quote:
 

 1) True. My bad. In my elog entry (but not in my fit code) I forgot the impedance Z= 750Ohm (as in the fit) of the resonant circuit in front of the square root: Vn = Vdn + Z * sqrt( 2 e Idc )

2) That is exactly the point I was raising! The measured dark noise at resonance is 2x what I expect.

3) I don't have uncertainties for the fit offset (that is, for the Dark Noise). The quick fit that I used (Matlab's Non Linear Least Squares method) doesn't provide 95% confidence bounds when I constrain the offset parameter the way I did (I forced it to be strictly positive).
Sure. It's not a very good fit. I just wanted to see how the data was going.

I also admitted that the data points were few, especially around the 3dB point.

Today I'm going to repeat the measurement with a new setup that lets me tune the light intensity more finely.

 Here's another measurement of the noise of the REFL11 PD.

This time I made the fit constraining the Dark Noise. I realized that it didn't make much sense leaving it as a free coefficient: the dark noise is what it is.

2010-04-09_REFL11NoiseMeasurements.png

Result: the transimpedance of REFL11at 11 MHz is about 4000 Ohm.

Note:
This time, more properly, I refer to the transimpedance as the ratio between Vout @11Mhz / Photocurrent. In past entries I improperly called transimpedance the impedance of the circuit which resonates with the photodiode.
  2789   Mon Apr 12 16:20:05 2010 AlbertoConfiguration40m UpgradingREFL55 improved
During the commissioning of the AS55 PD, I learned how to get a much better rejection of the 11MHz modulation.
So I went back to REFL55 and I modified it using the same strategy. (Basically I added another notch to the circuit).
After a few days of continuous back and forth between modeling, measuring, soldering, tuning I got a much better transfer function.

All the details and data will be included in the wiki page (and so also the results for AS55). Here I just show the comparison of the transfer functions that I measured and that I modeled.

I applied an approximate calibration to the data so that all the measurements would refer to the transfer function of Vout / PD Photocurrent. Here's how they look like. (also the calibration will be explained in the wiki)

2010-04-12_REFL55_TF_model_to_meas_comparison.png.

The ratio between the amplitude of the 55Mhz modulation over the 11MHz is ~ 90dB

The electronics TF doesn't provide a faithful reproduction of the optical response.

  2788   Mon Apr 12 14:20:10 2010 kiwamuUpdateGreen LockingPZT response for the innolight

I measured a jitter modulation caused by injection of a signal into laser PZTs.

The measurement has been done by putting a razor blade in the middle way of the beam path to cut the half of the beam spot, so that a change of intensity at a photodetector represents the spatial jitter of the beam.

However the transfer function looked almost the same as that of amplitude modulation which had been taken by Mott (see the entry).

This means the data is dominated by the amplitude modulation instead of the jitter. So I gave up evaluating the data of the jitter measurement.

  2787   Sun Apr 11 19:05:34 2010 KojiOmnistructureComputersWhere are the laptops?

One dell is in the clean room for the suspension work.

Quote:

I can't find the DELL laptop anywhere in the lab. Does anyone know where it is?

Also one of the two netbooks is missing.

 

  2786   Sun Apr 11 13:51:04 2010 AlbertoOmnistructureComputersWhere are the laptops?

I can't find the DELL laptop anywhere in the lab. Does anyone know where it is?

Also one of the two netbooks is missing.

  2785   Fri Apr 9 06:45:28 2010 AlbertoUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Quote:

Something must be wrong. 

1. Physical Unit is wrong for the second term of "Vn = Vdn + Sqrt(2 e Idc)"

2. Why does the fit go below the dark noise?

3. "Dark noise 4 +/- NaN nV/rtHz"   I can not accept this fitting.

Also apparently the data points are not enough.

 1) True. My bad. In my elog entry (but not in my fit code) I forgot the impedance Z= 750Ohm (as in the fit) of the resonant circuit in front of the square root: Vn = Vdn + Z * sqrt( 2 e Idc )

2) That is exactly the point I was raising! The measured dark noise at resonance is 2x what I expect.

3) I don't have uncertainties for the fit offset (that is, for the Dark Noise). The quick fit that I used (Matlab's Non Linear Least Squares method) doesn't provide 95% confidence bounds when I constrain the offset parameter the way I did (I forced it to be strictly positive).
Sure. It's not a very good fit. I just wanted to see how the data was going.

I also admitted that the data points were few, especially around the 3dB point.

Today I'm going to repeat the measurement with a new setup that lets me tune the light intensity more finely.

  2784   Thu Apr 8 20:53:13 2010 KojiUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Something must be wrong. 

1. Physical Unit is wrong for the second term of "Vn = Vdn + Sqrt(2 e Idc)"

2. Why does the fit go below the dark noise?

3. "Dark noise 4 +/- NaN nV/rtHz"   I can not accept this fitting.

Also apparently the data points are not enough.

Quote:

From the measurements of the 11 MHz RFPD at 11Mhz I estimated a transimpedance of about 750 Ohms. (See attached plot.)

The fit shown in the plot is: Vn = Vdn + sqrt(2*e*Idc) ; Vn=noise; Vdn=darknoise; e=electron charge; Idc=dc photocurrent

The estimate from the fit is 3-4 times off from my analsys of the circuit and from any LISO simulation. Likely at RF the contributions of the parassitic components of each element make a big difference. I'm going to improve the LISO model to account for that.

2010_04_05_REFL11_ShotnoiseVsPhotocurrent.png

The problem of the factor of 2 in the data turned out to be not a real one. Assuming that the dark noise at resonance is just Johnson's noise from the resonant circuit transimpedance underestimates the dark noise by 100%.

 

  2783   Thu Apr 8 10:24:33 2010 AlbertoUpdate40m UpgradingREFL11 Noise Vs Photocurrent

Quote:

From the measurements of the 11 MHz RFPD at 11Mhz I estimated a transimpedance of about 750 Ohms. (See attached plot.)

Putting my hands ahead, I know I could have taken more measurements around the 3dB point, but the 40m needs the PDs soon.

  2782   Thu Apr 8 10:17:52 2010 AlbertoUpdate40m UpgradingREFL11 Noise Vs Photocurrent

From the measurements of the 11 MHz RFPD at 11Mhz I estimated a transimpedance of about 750 Ohms. (See attached plot.)

The fit shown in the plot is: Vn = Vdn + sqrt(2*e*Idc) ; Vn=noise; Vdn=darknoise; e=electron charge; Idc=dc photocurrent

The estimate from the fit is 3-4 times off from my analsys of the circuit and from any LISO simulation. Likely at RF the contributions of the parassitic components of each element make a big difference. I'm going to improve the LISO model to account for that.

2010_04_05_REFL11_ShotnoiseVsPhotocurrent.png

The problem of the factor of 2 in the data turned out to be not a real one. Assuming that the dark noise at resonance is just Johnson's noise from the resonant circuit transimpedance underestimates the dark noise by 100%.

  2781   Wed Apr 7 11:11:19 2010 AlbertoUpdateElectronicsREFL11 Noise Simulation

Quote:

What??? I don't see any gray trace of Rs in the plot. What are you talking about?

Anyway, if you are true, the circuit is bad as the noise should only be dominated by the thermal noise of the resonant circuit.

Quote:
LISO simulations confirm the estimate of ~15nV for the noise of REFL11.
The largest contribution comes from the output resistor (Rs in the schematic below).
See attached plot.

 

The colors in the plot were misleading.
Here's hopefully a better plot.
The dominant sources of noise are the resonant of the photodiode (~10Ohm), the max4107, the resistor in series to ground at the - input of the max4107.
Attachment 1: rfpd11_testinput_noiseplot.pdf
rfpd11_testinput_noiseplot.pdf
  2780   Wed Apr 7 10:58:15 2010 KojiUpdateElectronicsREFL11 Noise Simulation

What??? I don't see any gray trace of Rs in the plot. What are you talking about?

Anyway, if you are true, the circuit is bad as the noise should only be dominated by the thermal noise of the resonant circuit.

Quote:
LISO simulations confirm the estimate of ~15nV for the noise of REFL11.
The largest contribution comes from the output resistor (Rs in the schematic below).
See attached plot.

 

  2779   Wed Apr 7 10:48:04 2010 AlbertoUpdateElectronicsREFL11 Noise Simulation
LISO simulations confirm the estimate of ~15nV for the noise of REFL11.
The largest contribution comes from the 50Ohm output resistor (Rs in the schematic below), the 450Ohm feedback resistor of the max4107 opamp stage; the 10KOhm resistor at the Test Input connector.
 
See attached plot.
 
(It's also all in the SVN, under https://nodus.ligo.caltech.edu:30889/svn/trunk/alberto/40mUpgrade/RFsystem/RFPDs/)
#
#                 gnd
#                 |
#                 Cw2
#                 |
#                 n23
#                 |
#                 Lw2
#                 |
#   gnd           n22
#   |             |
#   Rip           Rw2
#   |             |                   |\
#   nt- Rsi-n2- - - C2 - n3 -  - -  - |  \
#            |    |      |   |        |4106>-- n5 - Rs -- no                                                            
# iinput    Rd   L1     L2 R24    n6- |  /     |           |
#    |- nin- |    |      |   |    |   |/       |         Rload
#           Cd   n7     R22 gnd   |            |           |
#            |    |      |        | - - - R8 - -          gnd
#           gnd  R1     gnd      R7
#                 |               |
#                gnd             gnd
#
#
#
Attachment 1: rfpd11_testinput_noiseplot.pdf
rfpd11_testinput_noiseplot.pdf
  2778   Wed Apr 7 09:00:01 2010 steveHowToPEMprepare to open chamber

In order to minimize the diffusion of more dust particles into the vented IFO vacuum envelope

BEFORE opening chamber:

-Have a  known plan,

-Heavy 1" thick door requires 3 persons- of  one experienced and one certified crane operator and steel tow safety shoes

-Block IFO beams, be ware of experimental set up of other hazards: 1064,  visible or new-special installation

- Look at the particle counter, do not open above 6,000 particles of 0.5 micron. Construction activities are winding down. See  plot of 35 days since we  vented.

-Have clean door stand for heavy door, covered with merostate at the right location and dry-clean screws for light covers,

-Prepare lint free wipers for o-rings,(no solvent on o-ring!) Kimwipes for outside of chamber and metal covers, methanol and powder free gloves

-Wipe with wet Kimwipe-tissue of methanol around the door, chamber of interest and o-ring cover ring

-Cut door covering merostate and tape it into position,..if in place...check  folded-merostate position, if dusty... replace it

-Is your cleanroom garment clean?.......if in doubt ....replace it

-Keep surrounding area free and clean

-Make sure that HEPAs are running: PSL-enclosure, two mobile units and south end flow banch

-Check the tools: are they really clean? wipe it with wet Kimwipe, do you see anything on the Kimwipe?

 

-You are responsible to close chamber ASAP with light door or doors as you finished for the day.

Merostate cover down is appropriate during daily brakes.

Attachment 1: 0.5micron.jpg
0.5micron.jpg
  2777   Tue Apr 6 22:54:34 2010 KojiUpdateSUSITMY (south) aligned

Kiwamu and Koji

ITMY (south) was aligned with regard to the 40m-long oplev with the green laser pointer. Now the cavity is waiting for the green light injected from the end table

The OSEMs were adjusted with the aligned optics, but still a bit off from the center. They need to be adjusted again.
One round-shaped counter-weight removed from the table. Some counter weights are moved.

Some tools and the level gauge were removed from the table.

BAD news: I could clearly see scatter of the green beam path because of the dusts in the arm tube. Also many dusts are seen on the ITM surface.

 

Picture of the ETM - reflection from the ITM is hitting the mirror and the suspension structures.

IMG_2362.jpg

 


1. Shoot the ITM center with the green beam.

- Two persons with walkie-talkies required for this work.

- Turn on the end green pointer. We could see the long trace of the beam sliced by the beam tube wall.

- Look at the tube peeping mirror for the CCD.

- Adjust yaw such that the beam trace on the tube wall is parallel to the arm.

- Adjust pitch such that the beam trace on the tube gets longer. This means that spot gets closer to the ITM.

- Continue pitch adjustment until some scatter appears on the ITM tower.

- Once the spot appears on the tower, you can easily adjust it on the mirror

2. Adjust pitch/yaw bias such that the reflection hits the ETM.

- Initially the ITM alignment is totally bad. ==> You clealy see the spot on the wall somewhere close to the ITM.

- Adjust pitch/yaw bias such that the spot goes farther as far as possible.

- Once you hit the suspension tower, the scatter is obviously seen from the peeping mirror.

- You can match the incident beam and the scattering of the reflection. You also can see the reflection from the ETM towards the ITM as the spot size gets huge (1/2 tube diameter).

- We found that the bias is ~-2 for pitch and ~-6 for yaw.

3. Go into the chamber. Check the table leveling.

- Open the light door.

- I found that the table is not leveled. Probably it drifted after the move of the weight (i.e. MOS removal).

- Removed one of the round-shaped weight. Moved the other weights such that the table was leveled.

4. Remove the bias for yaw and rotate suspension tower such that the reflection hit the center of the ETM.

- Removed the yaw bias. This makes the reflected spot totally off from the ETM.

- Rotate suspension tower so that the beam can approximately hit the ETM.

- Look at the peeping mirror, the beam is aligned to the ETM.

5. Adjust OSEMs

- Push/pull the OSEMs such that we have the OSEM outputs at the half of the full scale.

6. Adjust alignment by the bias again.

- Moving OSEMs changes the alignment. The pitch/yaw biases were adjusted to have the beam hitting on the ETM.

- Bias values at  the end of the work: Pitch -0.8159 / Yaw -1.2600

7. Close up the chamber

- Remove the tools and the level gauge.

- Close the light door.

  2776   Tue Apr 6 16:55:28 2010 AlbertoUpdateComputer Scripts / ProgramsData formats in the Agilent AG4395a Spectrum Analyzer

Quote:

Lately I've been trying to sort out the problem of the discrepancy that I noticed between the values read on the spectrum analyzer's display and what we get with the GPIB interface.

It turns out that the discrepancy originates from the two data vector that the display and the GPIB interface acquire. Whereas the display shows data in "RAW" format, the GPIB interface, for the way the netgpibdata script is written, acquires the so called "error-corrected data". That is the GPIB downloaded data is postprocessed and corrected for some internal calibration factors of the instrument.

Another problem that I noticed in the GPIB downloaded data when I was measuring noise spectrum, is an unwanted factor of 2 in the amplitude spectral density.
For example, measuring the amplitude spectral density of the FSS RF PD's dark noise at its resonant frequency (~21.5 MHz), I would expect ~15nV/rtHz from the thermal noise - as Rana pointed out in the elog entry 2759). However, the spectrum analyzer reads 30nV/rtHz, in both the display and the GPIB downloaded data, except for the above mentioned little discrepancy between the two. (The discrepancy is about 0.5dBm/Hz in the power spectrum density).
 
My measurement, as I showed it in the elog entry 2760) is of ~15nV/rtHz, but only becasue I divided by 2. Now I realize that that division was unjustified.
 
I'm trying to figure out the reason for that. By now I'm not sure we can trust the netgpib package for spectrum measurements with the AG4395.

 I noticed that someone, that wasn't me, has edited the wiki page about the netgpibdata under my name saying:

 " [...]

* A4395 Spectrum Units
Independetly by which unites are displayed by the A4395 spectrum analyzer on the screen, the data is saved in Watts/rtHz
"

That is not correct. The spectrum is just in Watts, since it gives the power over the bandwidth. The correspondent power spectral density is showed under the "Noise" measurement format and it's in Watts/Hz.
Watts/rtHz is not a correct unit.
  2775   Tue Apr 6 11:27:11 2010 AlbertoUpdateComputer Scripts / ProgramsData formats in the Agilent AG4395a Spectrum Analyzer

Lately I've been trying to sort out the problem of the discrepancy that I noticed between the values read on the spectrum analyzer's display and what we get with the GPIB interface.

It turns out that the discrepancy originates from the two data vector that the display and the GPIB interface acquire. Whereas the display shows data in "RAW" format, the GPIB interface, for the way the netgpibdata script is written, acquires the so called "error-corrected data". That is the GPIB downloaded data is postprocessed and corrected for some internal calibration factors of the instrument.

Another problem that I noticed in the GPIB downloaded data when I was measuring noise spectrum, is an unwanted factor of 2 in the amplitude spectral density.
For example, measuring the amplitude spectral density of the FSS RF PD's dark noise at its resonant frequency (~21.5 MHz), I would expect ~15nV/rtHz from the thermal noise - as Rana pointed out in the elog entry 2759). However, the spectrum analyzer reads 30nV/rtHz, in both the display and the GPIB downloaded data, except for the above mentioned little discrepancy between the two. (The discrepancy is about 0.5dBm/Hz in the power spectrum density).
 
My measurement, as I showed it in the elog entry 2760) is of ~15nV/rtHz, but only becasue I divided by 2. Now I realize that that division was unjustified.
 
I'm trying to figure out the reason for that. By now I'm not sure we can trust the netgpib package for spectrum measurements with the AG4395.
  2773   Mon Apr 5 14:10:06 2010 steveUpdateSUSsus damping restored

Quote:

This morning, at about 12 Koji found all the front-ends down.

At 1:45pm rebooted ISCEX, ISCEY, SOSVME, SUSVME1, SUSVME2, LSC, ASC, ISCAUX

Then I burtestored ISCEX, ISCEY, ISCAUX to April 2nd, 23:07.

The front-ends are now up and running again.

 I restored damping to all SUSes except ITM-east. The ITMX OSEMs are being used in the clean assembly room.

  2772   Mon Apr 5 13:52:45 2010 AlbertoUpdateComputersFront-ends down. Rebooted

This morning, at about 12 Koji found all the front-ends down.

At 1:45pm rebooted ISCEX, ISCEY, SOSVME, SUSVME1, SUSVME2, LSC, ASC, ISCAUX

Then I burtestored ISCEX, ISCEY, ISCAUX to April 2nd, 23:07.

The front-ends are now up and running again.

  2771   Mon Apr 5 13:20:16 2010 KojiOmnistructureElectronicssoldering iron broken

Albeto and Koji

We took the tip replacement from the blue tower.

I am looking at http://www.cooperhandtools.com/brands/weller/ for ordering the tips.

The burnt one seems to be "0054460699: RT6 Round Sloped Tip Cartridge for WMRP Pencil" We will buy one.

The replaced one is "0054460299: RT2 Fine Point Cartridge for WMRP Pencil" We will buy two.

I like to try this: "0054460999: RT9 Chisel Tip Cartridge for WMRP Pencil" We will buy one.

Quote:

This morning the pencil soldering iron of our Weller WD2000M Soldering Station suddenly stopped working and got cold after I turned the station on. The unit's display is showing a message that says "TIP". i checked out the manual, but it doesn't say anything about that. I don't know what it means. Perhaps burned tip?

Before asking Steve to buy a new one, I emailed Weller about the problem.

 

Attachment 1: weller_tips.jpg
weller_tips.jpg
  2770   Mon Apr 5 13:07:36 2010 JenneOmnistructureElectronicssoldering iron broken

Quote:

This morning the pencil soldering iron of our Weller WD2000M Soldering Station suddenly stopped working and got cold after I turned the station on. The unit's display is showing a message that says "TIP". i checked out the manual, but it doesn't say anything about that. I don't know what it means. Perhaps burned tip?

Before asking Steve to buy a new one, I emailed Weller about the problem.

 There should be a supply of extra tips in the Blue Spinny Cabiney (I can never remember it's French name....)  The drawer is something like the top row of one of the bottom sets of drawers.  You can pick the shape of tip you want, and stick it in.

  2769   Mon Apr 5 11:39:41 2010 steveUpdateSUSITM-south installation

Quote:

Steve and Koji (Friday, Apr 02)

Summary

Installation of ITMs are going on. Two new ITMs were placed on the optical table in the vacuum chambers. ITM for the south arm was put at the right place in accordance to the CAD drawing. ITM for the east arm is still at a temporaly place.


Tower placement (10:30-11:30)

- Put the tower on the table at a temporary place such that we can easily work on the OSEMs.

ITM (South arm) (14:00-16:30)

- Put the tower on the table at a temporary place such that we can easily work on the OSEMs.

- Leveled the table approximately.

- Released the EQ stops

- Removed anchors for the OSEM cables as it was too short. The wire distribution will be changed later.

- Put the OSEMs. Adjust the insertion to the middle of the OSEM ranges.

- Clamped the EQ stops again

- Placed the tower to the right place according to the CAD drawing.

- Released the EQ stops again.

- Check the OSEM values. The LL sensor showed small value (~0.5). Needs to be adjusted.

 


ITM (South) damping adjustment

- Found the signs for the facing magnets are reversed.

- Otherwise it damps very well.

 

 The cabling on the seismic stack was rerouted so it could reach the south edge of the table: the cables  were removed from the viton padded clamps and repositioned this morning.

ITM-south tower's earthquake screw viton tips could be a little bit larger. They do not stay in their screw hole after a hard clamping action.

4-40 earthquake screws under the test mass:viton tips can fall out without action, the treads are cross threaded so the screws are wobbling

 

  2768   Mon Apr 5 10:33:12 2010 AlbertoOmnistructureElectronicssoldering iron broken

This morning the pencil soldering iron of our Weller WD2000M Soldering Station suddenly stopped working and got cold after I turned the station on. The unit's display is showing a message that says "TIP". i checked out the manual, but it doesn't say anything about that. I don't know what it means. Perhaps burned tip?

Before asking Steve to buy a new one, I emailed Weller about the problem.

  2767   Mon Apr 5 10:23:40 2010 AlbertoUpdate40m UpgradingREFL11 Low Frequency Oscilaltion Reduced

After adding an inductor L=100uH and a resistor R=10Ohm in parallel after the OP547A opamp that provide the bias for the photodiode of REFL11, the noise at low frequency that I had observed, was significantly reduced.

See this plot:

 2010_04_05_REFL11_darknoise_with_100uH_coil_10ohm_res.png

A closer inspection of the should at 11MHz in the noise spectrum, showed some harmonics on it, spaced with about 200KHz. Closing the RF cage and the box lid made them disappear. See next plot:

 2010_04_05_REFL11_darknoise_wide_freq_window_lid_open-closed.png

The full noise spectrum looks like this:

2010_04_05_REFL11_darknoise_wide_freq_window_lid_open-closed.png

A big bump is present at ~275MHz. it could important if it also shows up on the shot noise spectrum.

  2766   Mon Apr 5 09:48:57 2010 KojiUpdateSUSITMs placed on the tables in the chambers

Steve and Koji (Friday, Apr 02)

Summary

Intsallation of ITMs are going on. Two new ITMs were placed on the optical table in the vacuum chambers. ITM for the south arm was put at the right place in accordance to the CAD drawing. ITM for the east arm is still at a temporaly place.


Tower placement (10:30-11:30)

- Put the tower on the table at a temporary place such that we can easily work on the OSEMs.

ITM (South arm) (14:00-16:30)

- Put the tower on the table at a temporary place such that we can easily work on the OSEMs.

- Leveled the table approximately.

- Released the EQ stops

- Removed anchors for the OSEM cables as it was too short. The wire distribution will be changed later.

- Put the OSEMs. Adjust the insertion to the middle of the OSEM ranges.

- Clamped the EQ stops again

- Placed the tower to the right place according to the CAD drawing.

- Released the EQ stops again.

- Check the OSEM values. The LL sensor showed small value (~0.5). Needs to be adjusted.

 


ITM (South) damping adjustment

- Found the signs for the facing magnets are reversed.

- Otherwise it damps very well.

 

  2765   Mon Apr 5 08:43:48 2010 steveUpdatePEMearthquake mag 7.2

Large earthquake shakes Baja California, Mexico and 6 over Magnitude 5 aftershakes follow.  The frontend computers are still down since Friday.

Attachment 1: eq7.2.jpg
eq7.2.jpg
  2764   Mon Apr 5 01:02:07 2010 ranaUpdatePEMGuralp interface box turned off

I was checking into the Guralp situation today. I put the rubber balls underneath the granite block (the Q is too high), but found unfortunately that Jenne's styrofoam box is too short to cover the Guralps on top of the granite. If the box was skinny enough to fit on the block or taller by ~6 inches, it would be perfect. We need some new Seismo boxes.

 

Here's the story of the Gur2 noise so far. We need to pull out and repair the breakout box.

1) At some point we noticed that the Guralp2 X channel was behaving badly.

2) Steve tried recentering with just a +12V supply - this didn't work. Jenne then centered it using the +/- 12V supply. This was OK.

3) Around noon on March 24, the channel 'goes bad' again.

4) On the afternoon of the 25th, most of the channels go to zero, but the GUR2X channel stays bad. There's NO ENTRY in the elog about this. This is UNACCEPTABLE. Apparently, the seismometers were disconnected without shutting off the power to the box. You MUST elog everything - otherwise, go home and sit on your hands.

5) On the evening of the 31st, Steve turns off the Guralp breakout box. From the trend, you can see that the signals all go to zero at that time.

6) From then until today, there is no noise in the GUR2X channel. From these tests we can guess that the problem is in the GUR2X channel of the breakout box, but not in the AA Chassis or the ADC, since those showed no excess noise with the box turned OFF. Its hard to be sure without elog entries, but I assume that 3/25-3/31 was a 'seismometer disconnected', but 'box on' state.

Attachment 1: Untitled.png
Untitled.png
  2763   Sun Apr 4 17:32:07 2010 AlbertoMetaphysicsGeneralnew y-arm?

Quote:

There's several more of the this vintage in one of the last cabinets down the new Y-arm.

 Hold on, did the arms get re-baptized?

  2762   Sun Apr 4 00:21:42 2010 rana, kojiSummaryElectronicsCheckout of EG&G (PARC) preamp model #113, s/n 49135

We tested out the functionality of the EG&G 113 preamp that I found in one of the cabinets. This is one of the ancestors of the SR560 preamp that we are all used to.

It turns out that it works just fine (in fact, its better than the SR560). The noise is below 3nV/rHz everywhere above 30 Hz. The filter settings from the front panel all seem to work well. And the red knob on the front panel allows for continuous (i.e. not steps) gain adjustment. In the high-bandwidth mode (low pass filter at 300 kHz), there is ~35 deg of phase lag at 100 kHz. So the box is pretty fast.

IMG_0628.JPG

I would easily recommend this above the SR560 for use in all applications where you don't need to drive a 50 Ohm load. Also the battery is still working after 17 years!

There's several more of the this vintage in one of the last cabinets down the new Y-arm.

  2761   Sat Apr 3 19:54:19 2010 AlbertoUpdate40m UpgradingREFL11 and REFL55 PDs Noise Spectrum

These are the dark noise spectrum that I measured on the 11MHz and 55MHz PD prototypes I modified.

The plots take into account the 50Ohm input impedance of the spectrum analyzer (that is, the nosie is divided by 2).

2010-04-03_REFL11_darknoise.png 2010-04-03_REFL55_darknoise.png

With an estimated transimpedance of about 300Ohm, I would expect to have 2-3nV/rtHz at all frequencies except for the resonant frequencies of each PD. At those resonances I would expect to have ~15nV/rtHz (cfr elog entry 2760).

Problems:

  1. For the 55MHz PD the resonance peak is too small
  2. In the 55 MHz: noise is present at about 7MHz
  3. In the 11MHz PD there's a lot of noise below 10 MHz.

I have to figure out what are the sources of such noises.

Suggestions?

  2760   Sat Apr 3 16:07:40 2010 AlbertoConfigurationPSLReference Cavity PD Noise Spectrum

 I was aware of a problem on those units since I acquired the data. Then it wasn't totally clear to me which were the units of the data as downloaded from the Agilent 4395A, and, in part, still isn't.

It's clear that the data was in units of spectrum, an not spectral density: in between the two there is a division by the bandwidth (100KHz, in this case). Correcting for that, one gets the following plot for the FSS PD:

2010-03-29_FSS_PD_shotnoise_and_darknoise.png

Although the reason why I was hesitating to elog this other plot is that it looks like there's still a discrepancy of about 0.5dBm between what one reads on the display of the spectrum analyzer and the data values downloaded from it.

However I well know that, I should have just posted it, including my reserves about that possible offset (as I'm doing now).

Quote:

The units on this plot are completely bogus - we know that the thermal noise from the resonant part of the circuit is just V = sqrt(4*k*T*Z) ~ 3nV/rHz. Then the gain of the MAX4107 stage is 10. The output resistor is 50 Ohms, which forms a divide by 2 with the input impedance of the spectrum analyzer and so the bump in the dark noise should only be 15 nV/rHz and not microVolts.

Quote:

[Rana, Alberto]

This evening we measured the noise spectrum of the reference cavity PD used in the FSS loop. From that we estimated the transimpedance and found that the PD is shot-noise limited. We also found a big AM oscillation in correspondence of the FSS modulation sideband which we later attenuated at least in part.

This plot shows the spectrum noise from the RF output of the photodetector.

  2759   Sat Apr 3 11:35:47 2010 ranaConfigurationPSLReference Cavity PD Noise Spectrum

The units on this plot are completely bogus - we know that the thermal noise from the resonant part of the circuit is just V = sqrt(4*k*T*Z) ~ 3nV/rHz. Then the gain of the MAX4107 stage is 10. The output resistor is 50 Ohms, which forms a divide by 2 with the input impedance of the spectrum analyzer and so the bump in the dark noise should only be 15 nV/rHz and not microVolts.

Quote:

[Rana, Alberto]

This evening we measured the noise spectrum of the reference cavity PD used in the FSS loop. From that we estimated the transimpedance and found that the PD is shot-noise limited. We also found a big AM oscillation in correspondence of the FSS modulation sideband which we later attenuated at least in part.

This plot shows the spectrum noise from the RF output of the photodetector.

  2758   Fri Apr 2 08:52:21 2010 AlbertoUpdateelogelog restarted

i just restarted the elog for the third time in the past 12 hours.

I checked the elog.log file to debug the problem. It doesn't contain eveidence of any particular cause, except for png/jpg file uploads happened last night.

I'm not sure we can blame Image Magic again because the last crash seems to be occurred just after an entry with e jpg picture was included in the body of the message. I think Image Magic is used only for previews of attachments like pdfs or ps.

Maybe we should totally disable image magic.

  2757   Thu Apr 1 20:29:02 2010 HartmutUpdateGreen Lockingsimple PD test circuit

I made a simple PD test circuit which may allow to test PD response up to few 100MHz.

Its not for low noise, only for characterising PD response.

Here is the circuit:

The 2 capacitor values (for bypassing) are kind of arbitrary, just what I found around

(one medium, one small capacity). Could be improved by better RF types (e.g. Mica).

The PD type has no meaning. I put in the Centronic 15-T5 for a start.

The bias can be up to 20V for this diode.

The signal appears across R1. It is small, to make a large bandwidth.

R2 is just for slightly decoupling the signal from the following RF amplifier.

The wire into the RF amplifier is short (~cm). And the amplifier is supposed to have 50 Ohm

input impedance.

I use a mini circuits ZFL 500 here.

power supply for this is 15V.

pdtest.png

  2756   Thu Apr 1 19:59:32 2010 MottUpdateGeneralPZT response for the innolight

 

 We measured the Amplitude Modulation response of the PZTs, to find regions with large phase modulation but small amplitude modulation.

We did this by blocking 1 arm of the PLL, feeding the source output of the Network Analyzer into the PZT input of the laser in question, and reading the output of the PD on the NA.  We calibrated by dividing by the DC voltage of the PD (scaled by the ratio of the AC gain to DC gain of the New Focus PD).

The AM response of the Innolight looks fairly smooth up to ~1MHz, and it is significantly below the PM response for most of its range.  The region between 20 and 30 kHz shows very good separation of about 10^3 rad/RIN (and up to 10^5 rad/RIN at ~21.88 kHz, where there is the negative spike in the AM response). The region between 1.5 MHz and 2MHz also looks viable if it is desirable to actuate at higher frequencies.

The Lightwave offers very good AM/PM separation up to about 500 kHz, but becomes quite noisy about 1MHz.

Attachment 1: Innolight_AM_Response.png
Innolight_AM_Response.png
Attachment 2: Innolight_AM_PM.png
Innolight_AM_PM.png
Attachment 3: InnoVsLW_PM.png
InnoVsLW_PM.png
Attachment 4: Innolight_AM_Response.png
Innolight_AM_Response.png
Attachment 5: Lightwave_AM_PM.png
Lightwave_AM_PM.png
  2755   Thu Apr 1 18:44:40 2010 KojiUpdateGeneralPZT response for the innolight

Innolight 10 rad/V @ 100kHz => 1e6/f rad/V => 1MHz/V

LWE 30 rad/V @ 100kHz => 3e6/f rad/V => 3MHz/V

---------

BTW, don't let me calculate the actuator response everytime.

The elog (=report) should be somewhat composed by the following sections

Motivation - Method - Result (raw results) - Discussion (of the results)

Quote:

  We realized that we had measured the wrong calibration value; we were using the free-running error signal with the marconi far from the beat frequency, which was very small.  When we put the Marconi right at the beat, the signal increased by a factor of ~12 (turning our original calibration of 10 mV/rad into 120 mV/rad).  The re-calibrated plots are attached. 

 

  2754   Thu Apr 1 18:05:29 2010 MottUpdateGeneralPZT response for the innolight

 

 We realized that we had measured the wrong calibration value; we were using the free-running error signal with the marconi far from the beat frequency, which was very small.  When we put the Marconi right at the beat, the signal increased by a factor of ~12 (turning our original calibration of 10 mV/rad into 120 mV/rad).  The re-calibrated plots are attached. 

Attachment 1: Innolight_Response_calFix.png
Innolight_Response_calFix.png
Attachment 2: Lightwave_response_calFix.png
Lightwave_response_calFix.png
  2753   Thu Apr 1 17:35:24 2010 KojiUpdateSUSWorking on ITMX/Y

Steve and Koji

- We removed old ITMX/Y from the chambers. Now they are temporarily placed on the flow table at the end. Steve is looking for nice storages for the 5inch optics.

- We wiped new ITMX/Y by isopropanol as they were dusty.

- We put them into the corresponding towers. Checked the balancing and magnet arrangements with the OSEMs. They were totally fine.

- We clamped the mirrors by the EQ stops. Wrapped the towers by Al foils.

Tomorrow we will put them into the chambers.

 

Attachment 1: IMG_2353.jpg
IMG_2353.jpg
  2752   Thu Apr 1 16:34:29 2010 HartmutUpdateGreen LockingSilicon PDs

just a few infos on Silicon PDs I looked up.

If you want to go beyond the 100MHz achievable with the device I worked on,

the one thing to improve is the opamp, where Steve is trying to find OPA657.

This is a FET with 1.6GHz BWP, minimum stable gain of 7, and 4.8nV/rt(Hz) noise.

Should be ok with 750-1000 Ohm transimpedance.

The other thing you might want to change is the PD

(although it might be the 1cm PD with high bias is as fast as smaller ones with lower bias).

There are two types of other Si diodes at the 40m right now (~3mm):

-Rana and I found a Centronic OSD 15-5T in the old equipment

-Frank gave me a Hamamatsu S1223-01 on a Thorlabs pre-amp device (could be taken out).

 

The Centronic OSD 15-5T has up to 80pF with 12 V bias according to the datasheet.

The Hamamatsu S1223-01 is stated with 20pF only, but stated to have a max. frequency resp. of 20MHz ('-3db point').

I dont know what this means, as the corner freq. of 10pF into 50Ohm is still 160MHz.

In any case there are faster 3mm types to start with, as for example Hamamatsu S3399 (~ 90$),

which is stated to have the corner at 100MHz with 50 Ohm load.

For this type the stated capacity (20pF) looks consistent with ~100MHz corner into 50 Ohm.

So probably you can get higher BW with this one using much smaller load, as in transimpedance stage.

 

 

  2751   Thu Apr 1 15:21:12 2010 ranaUpdateGreen Lockingfrequency counter for green PLL

 

  2750   Thu Apr 1 12:07:22 2010 ranaUpdateGeneralPZT response for the innolight

The Lightwave NPRO should be around 5 MHz/V. 

The Innolight PZT coefficient is ~1.1 MHz/V.

(both are from some Rick Savage LHO elog entries)

  2749   Thu Apr 1 10:47:48 2010 KojiUpdateGeneralPZT response for the innolight

Innolight: 100rad/V @ 100kHz  => 1e7/f rad/V => 10MHz/V

LWE: 500rad/V @ 100kHz =>  5e7/f rad/V => 50MHz/V

They sound little bit too big, aren't they?

  2748   Thu Apr 1 10:21:58 2010 MottUpdateGeneralPZT response for the innolight

Quote:

The shape of the TF looks nice but the calibration must be wrong.

Suppose 1/f slope with 10^-4 rad/V at 10kHz. i.e. m_pm = 1/f rad/V
This means m_fm = 1 Hz/V. This is 10^7 times smaller than that of LWE NPRO.

Quote:

Kiwamu and I measure the PZT response of the Innolight this evening from 24 kHz to 2MHz.  

We locked the PLL at ~50 MHz offset using the Lightwave NPRO and and swept the Innolight with the network analyzer (using the script I made; it has one peculiar property, but it does work correctly).  

We will post the plot of the Lightwave PZT response tomorrow morning.

 

 Koji is absolutely right.  I just double checked my matlab code, and saw that I divided when I should have multiplied.  The correctly calibrated plots are attached here for the Innolight and the lightwave.  Kiwamu and I will measure the amplitude and the jitter today.

Attachment 1: Innolight_Response.png
Innolight_Response.png
Attachment 2: Lightwave_response.png
Lightwave_response.png
ELOG V3.1.3-